OpenStudy (amtran_bus):

What does i^1, i^2, i^3, and i^4 equal in imaginary numbers?

OpenStudy (anonymous):

these are KEY facts that u should memorize i^1= sqrt(-1) i^2= -1 i^3= -sqrt(-1) i^4= 1

OpenStudy (amtran_bus):

Thanks God bless

OpenStudy (anonymous):

this also applies to multiples of these numbers

OpenStudy (amtran_bus):

SO it goes negative to negative to negative to positive?

OpenStudy (anonymous):

amazing

OpenStudy (amtran_bus):

I like school buses